OpenStudy (anonymous):

find an equation that has the given solutions: x=2i , x=-2i

OpenStudy (anonymous):

(x-2i)(x+2i)=0 x^2+2ix-2ix-4i^2=0 x^2 + 4 = 0

OpenStudy (anonymous):

thanks!